Calculate the length of the resistor of a circular wire used in a heater element connected to a 240V supply and the heater consumes 8 joules of energy in 0,25 seconds. The resistor element is made of aluminium and it has a uniform cross-sectional area with a diameter of 0,00016 inches.


Hints: 1inch = 25,4 mm, resistivity of aluminium = 2,825 x 10- 6 cm

1
Expert's answer
2022-03-21T12:45:41-0400

Power

"P=\frac{E}{t}"

"P=\frac{8}{0.25}=32W"

"P=\frac{V^2}{R}"

"32=\frac{240^2}{R}"

"R=\frac{240^2}{32}"

"R=1800\Omega"

Now we know that

"R=\frac{\rho l}{A}"

"l=\frac{RA}{\rho}"


"l=\frac{1800\times3.14\times(4.064\times10^{-6})^2}{2.825\times10^{-6}}=3.30\times10^{-5}m"